OpenStudy (anonymous):

complete the square x^2-6x-25=0

hero (hero):

Do you want the solution or a full explanation?

OpenStudy (anonymous):

either :)

hero (hero):

Well, I'll provide a solution. A full explanation is a bit much and if you don't want it, that makes this easier.

hero (hero):

x^2-6x - 25 = 0 x^2 -6x = 25 x^2-6x +9 = 25 + 9 (x-3)^2 = 34 x - 3 = sqrt{34} x = plus or minus sqrt(34) + 3

hero (hero):

That's just a solution. A full explanation would provide general steps on how to complete the square

OpenStudy (anonymous):

thanks this is perfect!